# Mitigating the Impact of Labeling Errors on Training via Rockafellian Relaxation

### v1

@@ Abstract
Labeling errors in datasets are common, if not systematic, in practice. They naturally arise in a variety of contexts—human labeling, noisy labeling, and weak labeling (i.e., image classification), for example. This presents a persistent and pervasive stress on machine learning practice. In particular, neural network (NN) architectures can withstand minor amounts of dataset imperfection with traditional countermeasures such as regularization, data augmentation, and batch normalization. However, major dataset imperfections often prove insurmountable. We propose and study the implementation of Rockafellian Relaxation (RR), a new loss reweighting, architecture-independent methodology, for neural network training. Experiments indicate RR can enhance standard neural network methods to achieve robust performance across classification tasks in computer vision and natural language processing (sentiment analysis). We find that RR can mitigate the effects of dataset corruption due to both (heavy) labeling error and/or adversarial perturbation, demonstrating effectiveness across a variety of data domains and machine learning tasks.
@@ 1 Introduction
Labeling errors are systematic in practice, stemming from various sources. For example, the reliability of human-generated labels can be negatively impacted by incomplete information, or the subjectivity of the labeling task - as is commonly seen in medical contexts, in which experts can often disagree on matters such as the location of electrocardiogram signal boundaries [8], prostate tumor region delineation, and tumor grading [20]. As well, labeling systems, such as Mechanical Turk [^1] http://mturk.com often find expert labelers being replaced with unreliable non-experts [27]. For all these reasons, it would be advisable for any practitioner to operate under the assumption that their dataset is corrupted with labeling errors, and possibly to a large degree.
In this paper, we propose a loss-reweighting methodology for the task of training a classifier on data having higher levels of labeling errors. We show that our method relates to optimistic and robust distributional optimization formulations aimed at addressing adversarial training (AT). These findings underscore our numerical experiments on NNs that suggest this method of training can provide test performance robust to high levels of labeling error, and to some extent, feature perturbation. Overall, we tackle the prevalent challenges of label corruption and class imbalance in training datasets, which are critical obstacles for deploying robust machine learning models. Our proposed approach implements Rockafellian Relaxations [23] to address corrupted labels and automatically manage class imbalances without the need for clean validation sets or sophisticated hyper-parameters - common constraints of current methodologies. This distinct capability represents our key contribution, making our approach more practical for handling large industrial datasets.
We proceed to discuss related works in section 2, and our specific contributions to the literature. In section 3 we discuss our methodology in detail and provide some theoretical justifications that motivate the effectiveness of our methodology. The datasets and NN model architectures upon which our experimental results are based are discussed in sections 4 and 5, respectively. We then conclude with numerical experiments and results in section 6.
@@ 2 Related Work
Corrupted datasets are of concern, as they potentially pose severe threats to classification performance of numerous machine-learning approaches [36], including, most notably, NNs [15, 33]. Naturally, there have been numerous efforts to mitigate this effect [28, 8]. These efforts can be categorized into robust architectures, robust regularization, robust loss function, loss adjustment, and sample selection [28]. Robust architecture methods focus on developing custom NN layers and dedicated NN architectures. This differs from our approach, which is architecture agnostic and could potentially "wrap around" these methods. While robust regularization methods like data augmentation [26], weight decay [16], dropout [29], and batch normalization [14] can help to bolster performance, they generally do so under lower levels of dataset corruption. Our approach, on the other hand, is capable of handling high levels of corruption, and can seamlessly incorporate methods such as these. In label corruption settings, it has been shown that loss functions, such as robust mean absolute error (MAE) [10] and generalized cross entropy (GCE) [35] are more robust than categorical cross entropy (CCE). Again, our method is not dependent on a particular loss function, and it is possible that arbitrary loss functions, including robust MAE and GCE, can be swapped into our methodology with ease. Our approach resembles the loss adjustment methods most closely, where the overall loss is adjusted based on a (re)weighting scheme applied to training examples.
@@
In loss adjustment methods, individual training example losses are typically adjusted multiple times throughout the training process prior to NN updates. These methods can be further grouped into loss correction, loss reweighting, label refurbishment, and meta-learning [28]. Our approach most closely resembles the loss reweighting methods. Under this scheme each training example is assigned a unique weight, where smaller weights are assigned to examples that have likely been corrupted. This reduces the influence of corrupted examples. A training example can be completely removed if its corresponding weight becomes zero. Indeed, a number of loss reweighting methods are similar to our approach. For example, Ren et al., [22] learn sample weights through the use of a noise-free validation set. Chang et al. [5] assign sample weights based on prediction variances, and Zhang et al. [34] examine the structural relationship among labels to assign sample weights. However, we view the need for a clean dataset, or at least one with sufficient class balance, by these methods as a shortcoming, and our method, in contrast, makes no assumption on the availability of such a dataset.
@@
Satoshi et al. [12] propose a two-phased approach to noise cleaning. The first phase trains a standard neural network to determine the top-$m$ most influential training instances that influence the decision boundary; these are subsequently removed from the training set to create a cleaner dataset. In the second phase, the neural network is retrained using the cleansed training set. Their method demonstrates superior validation accuracy for various values of $m$ on MNIST and CIFAR-10. Although impressive, their method does not address the fact that most industrial datasets have a reasonably large amount of label corruption [28] which, upon complete cleansing, could also remove informative examples that lie close to the decision boundary. Additionally, the value of $m$ is an additional hyper-parameter that could require significant tuning on different datasets and sources.
@@
Mengye et al. [22] propose dealing with label noise and class imbalance by learning exemplar weights automatically. They propose doing so in the following steps: a) Create a pristine noise-free validation set. b) Initially train on a large, noisy training dataset, compute the training loss on the training set, train on the clean validation set, and compute the training loss on the validation set. c) Finally, compute the exemplar weights that temper the training loss computed in step two with validation loss. This approach is algorithmically the most similar to ours, with some key differences. The major difference is that it treats noise and class imbalance similarly. Our approach deals with noisy labels explicitly and can cope with almost any amount of class imbalance automatically, as tested in our experiments with the open-source Hate-Speech dataset, where we experimented with different prevalence levels of Hate-Speech text. The biggest drawback of the method proposed by Mengye et al. is that it requires a clean validation set, which in practice is almost impossible to obtain; if it were possible, it would not be very prohibitive to clean the entire dataset. Noise, typically, is an artifact of the generative distribution which cannot be cherry-picked as easily in practice. Our approach does not require a clean dataset to be operational or effective.
@@ 3.1 Mislabeling
Let $\mathcal{X}$ denote a feature space, with $\mathcal{Y}$ a corresponding label space. Then $\mathcal{Z}:=\mathcal{X}\times\mathcal{Y}$ will be a collection of feature-label pairs, with an unknown probability distribution $D$. Throughout the forthcoming discussions, $\{(x_{i},y_{i})\}_{i=1}^{N}$ will denote a sample of $N$ feature-label pairs, for which some pairs will have a mislabeling. More precisely, we begin with a collection $(x_{i},\tilde{y}_{i})$ drawn i.i.d. from $D$, but there is some unknown set $C\subsetneq\{1,\ldots,N\}$ denoting (corrupted) indices for which $y_{i}=\tilde{y}_{i}$ if and only if $i\notin C.$ For those $i\in C,$ $y_{i}$ is some incorrect label, selected uniformly at random, following the Noise Completely at Random (NCAR) model [8] also known as uniform label noise.
@@ 3.2 Rockafellian Relaxation Method (RRM)
We adopt the empirical risk minimization (ERM) [31] problem formulation:
@@
In ERM it is common practice to assign each training observation $i$ a probability $p_{i}=1/N$. However, when given a corrupted dataset, we may desire to remove those samples that are affected; in other words, if $C\subsetneq\{1,...,N\}$ is the set of corrupted training observations, then we would desire to set the probabilities in the following alternative way:
$p=(p_{1},...,p_{N})\text{ with }p_{i}=\begin{cases}0,&\text{if }i\in C\\ \frac{1}{N-|C|},&\text{if }i\in\{1,...,N\}\setminus C,\end{cases}$ (2)
where $|C|$ is the cardinality of the unknown set $C$. In this work, we provide a procedure - the Rockafellian Relaxation Method (RRM) - with the intention of aligning the $p_{i}$ values closer to the desired (but unknown) $p$ of (2) in self-guided, automated fashion. It does so by adopting the Rockafellian Relaxation approach of [23]. More precisely, we consider the problem
$\min_{\theta}\Bigl{[}v(\theta):=\min_{u\in U}\sum_{i=1}^{N}(\frac{1}{N}+u_{i})\cdot J(\theta;x_{i},y_{i})+\gamma\lVert u\rVert_{1}\Bigr{]},$ (3)
where $U:=\{u\in\mathbb{R}^{N}:\sum_{i=1}^{N}u_{i}=0,\frac{1}{N}+u_{i}\geq 0\;\;\forall i=1,\ldots,N\}$, and some $\gamma>0$.
@@ 3.3 Analysis and Interpretation of Rockafellian Relaxation
Let $\gamma>0$ and $c=(c_{1},\dots,c_{N})\in\mathbb{R}^{N}$, with $c_{min}:=\min_{i}c_{i}$, and $c_{max}:=\max_{i}c_{i}$. Write $I_{min}:=\{i:c_{i}=c_{min}\}$, $I_{big}:=\{i:c_{i}=c_{min}+2\gamma\}$, and for any $S_{1}\subseteq I_{min},S_{2}\subseteq I_{big}$, define the polytope $U^{*}_{S_{1},S_{2}}:=\begin{Bmatrix}&u^{*}_{i}\geq 0\;\;\forall i:c_{i}=c_{min}\\ &u_{i}^{*}=0\;\;\forall i:c_{i}\in(c_{min},c_{min}+2\gamma)\\ u^{*}\in U:&u_{i}^{*}=-\frac{1}{N}\;\;\forall i\in I_{big}\setminus S_{2}\\ &u_{i}^{*}=-\frac{1}{N}\;\;\forall i:c_{i}>c_{min}+2\gamma\\ &u_{i}^{*}=0\;\;\forall i\in S_{1}\cup S_{2}\end{Bmatrix}.$ Then
$conv\left(\cup_{S_{1},S_{2}}U^{*}_{S_{1},S_{2}}\right)=\operatorname*{arg\,min}_{u\in U}\sum_{i=1}^{N}(\frac{1}{N}+u_{i})\cdot c_{i}+\gamma\lVert u\rVert_{1}.$ (4)
The theorem explains that the construction of any optimal solution $u^{*}$ essentially reduces to categorizing each of the losses among $\{c_{i}=J(\theta;x_{i},y_{i})\}_{i=1}^{N}$ as “small" or “big", according to their position in the partitioning of $[c_{min},\infty)=[c_{min},c_{min}+2\gamma)\cup[c_{min}+2\gamma,\infty)$. For losses that occur at the break points of $c_{min}$ and $c_{min}+2\gamma$, this classification can be arbitrary - hence, the use of $S_{1}$ and $S_{2}$ set configurations to capture this degree of freedom.
In particular, those points with losses $c_{i}$ exceeding $c_{min}+2\gamma$ are down-weighted to zero and effectively removed from the dataset. And in the event that $c_{max}-c_{min}<2\gamma$, no loss reweighting occurs. In this manner, while lasso produces sparse solutions in the model parameter space, RRM produces sparse weight vectors by assigning zero weight to data points with high losses.
Consequently, if $\chi:=\{i:c_{i}\in(c_{min}+2\gamma,\infty)\}$ converges over the course of any algorithmic scheme, e.g., Algorithm 1, to some set $C$, then we can conclude that these data points are effectively removed from the dataset even if the training of $\theta$ might proceed. This convergence was observed in the experiments of Section 6. It is hence of possible consideration to tune $\gamma$ for consistency with an estimate $\alpha\in[0,1]$ of labeling error in the dataset $\{(x_{i},y_{i})\}_{i=1}^{N}$. More precisely, we may tune $\gamma$ so that $\frac{|\chi|}{N}\approx\alpha.$

### v2

@@ Abstract
Labeling errors in datasets are common, arising in a variety of contexts, such as human labeling, noisy labeling, and weak labeling (i.e., image classification). Although neural networks (NNs) can tolerate modest amounts of these errors, their performance degrades substantially once error levels exceed a certain threshold. We propose a new loss reweighting, architecture-independent methodology, Rockafellian Relaxation Method (RRM) for neural network training. Experiments indicate RRM can enhance neural network methods to achieve robust performance across classification tasks in computer vision and natural language processing (sentiment analysis). We find that RRM can mitigate the effects of dataset contamination stemming from both (heavy) labeling error and/or adversarial perturbation, demonstrating effectiveness across a variety of data domains and machine learning tasks.
Machine Learning, ICML
@@ 1 Introduction
Labeling errors are systematic in practice, stemming from various sources. For example, the reliability of human-generated labels can be negatively impacted by incomplete information, or the subjectivity of the labeling task – as is commonly seen in medical contexts, in which experts can often disagree on matters such as the location of electrocardiogram signal boundaries (Frenay & Verleysen, 2014), prostate tumor region delineation, and tumor grading (Nir et al., 2018). As well, labeling systems, such as Mechanical Turk [^2] http://mturk.com often find expert labelers being replaced with unreliable non-experts (Snow et al., 2008). For all these reasons, it would be advisable for any practitioner to operate under the assumption that their dataset is contaminated with labeling errors, and possibly to a large degree.
In this paper, we propose a loss-reweighting methodology for the task of training a classifier on data having higher levels of labeling errors. More precisely, we provide a meta-algorithm that may “wrap” any loss-minimization methodology to add robustness. The method relates to optimistic and robust distributional optimization formulations aimed at addressing adversarial training (AT), which underscores our numerical experiments on NNs that suggest this method of training can provide test performance robust to high levels of labeling error, and to some extent, feature perturbation as well. Overall, we tackle the prevalent challenge of label contamination in training data, which is a critical obstacle for deploying robust machine learning models. The approach utilizes Rockafellian Relaxations (Royset et al., 2023) in addressing contaminated labels without the need for clean validation sets or sophisticated hyper-parameters - common constraints of current methodologies. This distinct capability represents a key contribution, making our approach more practical for handling large industrial datasets.
We proceed to discuss related works in Section 2, and our specific contributions to the literature. In Section 3 we discuss our methodology in detail and provide some theoretical justifications that motivate the effectiveness of our methodology. Sections 4 and 5 discuss the datasets and NN model architectures upon which our experimental results are based. We then conclude with numerical experiments and results in section 6.
@@ 2 Related Work
Contaminated datasets are of concern, as they potentially pose severe threats to classification performance of numerous machine-learning approaches (Zhu & Wu, 2004), including, most notably, NNs (Krause et al., 2016; Zhang et al., 2017). Naturally, there have been numerous efforts to mitigate this effect (Song et al., 2020; Frenay & Verleysen, 2014). These efforts can be categorized into robust architectures, robust regularization, robust loss function, loss adjustment, and sample selection (Song et al., 2020). Robust architecture methods focus on developing custom NN layers and dedicated NN architectures. This differs from our approach, which is architecture agnostic and could potentially ”wrap around” these methods. While robust regularization methods like data augmentation (Shorten & Khoshgoftaar, 2019), weight decay (Krogh & Hertz, 1991), dropout (Srivastava et al., 2014), and batch normalization (Ioffe & Szegedy, 2015) can help to bolster performance, they generally do so under lower levels of dataset contamination. Our approach, on the other hand, is capable of handling high levels of contamination, and can seamlessly incorporate methods such as these. In label contamination settings, it has been shown that loss functions, such as robust mean absolute error (MAE) (Ghosh et al., 2017), early learning regularization (ELR) (Liu et al., 2020), and generalized cross entropy (GCE) (Zhang & Sabuncu, 2018) are more robust than categorical cross entropy (CCE). Again, our method is not dependent on a particular loss function, and it is possible that arbitrary loss functions, including robust MAE and ELR, can be swapped into our methodology with ease. Our approach resembles the loss adjustment methods most closely, where the overall loss is adjusted based on a (re)weighting scheme applied to training examples.
@@
In loss adjustment methods, individual training example losses are typically adjusted multiple times throughout the training process prior to NN updates. These methods can be further grouped into loss correction, loss reweighting, label refurbishment, and meta-learning (Song et al., 2020). Our approach most closely resembles the loss reweighting methods. Under this scheme each training example is assigned a unique weight, where smaller weights are assigned to examples that have likely been contaminated. This reduces the influence of contaminated examples. A training example can be completely removed if its corresponding weight becomes zero. Indeed, a number of loss reweighting methods are similar to our approach. For example, Ren et al. (2018) learn sample weights through the use of a noise-free validation set. Chang et al. (2017) assign sample weights based on prediction variances, and Zhang et al. (2021) examine the structural relationship among labels to assign sample weights. However, we view the need by these methods for a clean dataset, or at least one with sufficient class balance, as a shortcoming, and our method, in contrast, makes no assumption on the availability of such a dataset.
@@
Hara et al. (2019) propose a two-phased approach to noise cleaning. The first phase trains a standard neural network to determine the top-$m$ most influential training instances that influence the decision boundary; these are subsequently removed from the training set to create a cleaner dataset. In the second phase, the neural network is retrained using the cleansed training set. Their method demonstrates superior validation accuracy for various values of $m$ on MNIST and CIFAR-10. Although impressive, their method does not address the fact that most industrial datasets have a reasonably large amount of label contamination (Song et al., 2020) which, upon complete cleansing, could also remove informative examples that lie close to the decision boundary. Additionally, the value of $m$ is an additional hyper-parameter that could require significant tuning on different datasets and sources.
@@
Perhaps the most similar approach to this work is that of Ren et al. (2018), in which label noise and class imbalance are treated simultaneously via the learning of exemplar weights. The biggest drawback of their method is that it requires a clean validation set, which in practice is almost impossible to obtain; if it were possible, it would not be very prohibitive to clean the entire dataset. Noise is typically an artifact of the generative distribution, which cannot be cherry-picked as easily in practice. Our approach does not require a clean dataset to be operational. Furthermore, even though we do not explicitly treat class imbalance, we are not affected by class imbalance, as demonstrated in our experiments with an open-source Toxic Comments dataset (Appendix Section C.1).
@@ 3.1 Mislabeling
Let $\mathcal{X}$ denote a feature space, with $\mathcal{Y}$ a corresponding label space. Then $\mathcal{Z}:=\mathcal{X}\times\mathcal{Y}$ will be a collection of feature-label pairs, with an unknown probability distribution $D$. Throughout the forthcoming discussions, $\{(x_{i},y_{i})\}_{i=1}^{N}$ will denote a sample of $N$ feature-label pairs, for which some pairs will be mislabeled. More precisely, we begin with a collection $(x_{i},\tilde{y}_{i})$ drawn i.i.d. from $D$, but there is some unknown set $C\subsetneq\{1,\ldots,N\}$ denoting (contaminated) indices for which $y_{i}=\tilde{y}_{i}$ if and only if $i\notin C.$ For indices $i\in C,$ $y_{i}$ is some incorrect label, say selected uniformly at random, following the Noise Completely at Random (NCAR) model (Frenay & Verleysen, 2014) also known as uniform label noise; however, the particular choice of noise/contamination matters not. While the experiments of Section 6 implement NCAR, we also consider non-uniform contamination in Appendix Section B.
@@ 3.2 Rockafellian Relaxation Method (RRM)
We adopt the empirical risk minimization (ERM) (Vapnik, 1991) problem formulation:
@@
In ERM it is common practice to assign each training observation $i$ a probability $p_{i}=1/N$. However, when given a contaminated dataset, we may desire to remove those samples that are affected; in other words, if $C\subsetneq\{1,...,N\}$ is the set of contaminated training observations, then we would desire to set the probabilities in the following alternative way:
$p=(p_{i})_{i\in[N]}\text{ with }p_{i}=\begin{cases}0,&\text{if }i\in C\\ \frac{1}{N-|C|},&\text{if }i\in[N]\setminus C,\end{cases}$ (2)
where $|C|$ is the cardinality of the unknown set $C$. In this work, we provide a procedure - the Rockafellian Relaxation Method (RRM) - with the intention of aligning the $p_{i}$ values closer to the desired (but unknown) $p$ of (2) in self-guided, automated fashion. It does so by adopting the Rockafellian Relaxation approach of (Royset et al., 2023). More precisely, with $\Delta(N)\subseteq\mathbb{R}^{N}$ denoting the standard probability simplex, and $p^{N}$ denoting the equally weighted (empirical) distribution $\{p^{N}_{i}\}_{i\in[N]}=\{1/N\}$, we consider the problem
$\min_{\theta}\min_{p\in\Delta(N)}\mathbb{E}_{(x,y)\sim p}\left[J(\theta;x,y)\right]+\gamma\cdot d_{TV}(p^{N},p)$
$=\min_{\theta}\Bigl{[}\underbrace{\min_{u\in U}\sum_{i=1}^{N}(\frac{1}{N}+u_{i})\cdot J(\theta;x_{i},y_{i})+\frac{\gamma}{2}\lVert u\rVert_{1}}_{v(\theta)}\Bigr{]},$ (3)
where $U:=\{u\in\mathbb{R}^{N}:\sum_{i=1}^{N}u_{i}=0,\frac{1}{N}+u_{i}\geq 0\;\;\forall i=1,\ldots,N\}$, $\gamma>0$, and $d_{TV}(p^{N},p):=\frac{1}{2}\sum_{i}|p^{N}_{i}-p_{i}|$. In words, in (3) an alternative distribution $p$ may now replace the empirical, at the cost of $\gamma$ per unit of total variation. Indeed, for any $u\in U$, we obtain a probability distribution $p^{u}\in\mathbb{R}^{N}$, defined via $p^{u}_{i}=1/N+u_{i},$ so that $\|u\|_{1}$ is the total variation distance between the empirical distribution and $p^{u}.$
@@ 3.3 Analysis and Interpretation of Rockafellian Relaxation
Let $\gamma>0$ and $c=(c_{1},\dots,c_{N})\in\mathbb{R}^{N}$, with $c_{min}:=\min_{i}c_{i}$, and $c_{max}:=\max_{i}c_{i}$. Write $I_{min}:=\{i:c_{i}=c_{min}\}$, $I_{mid}:=\{i:c_{i}\in(c_{min},c_{min}+\gamma)\}$, $I_{big}:=\{i:c_{i}=c_{min}+\gamma\}$, and for any $S_{1}\subseteq I_{min},$ $S_{2}\subseteq I_{big}$, define the polytope
$U^{*}_{S_{1},S_{2}}:=\begin{Bmatrix}&u^{*}\in U\\ &u^{*}_{i}\geq 0\;\;\forall i\in I_{min}\\ u^{*}:&u_{i}^{*}=0\;\;\forall i\in S_{1}\cup I_{mid}\\ &u^{*}_{i}\leq 0\;\;\forall i\in I_{big}\\ &u_{i}^{*}=-\frac{1}{N}\;\;\forall i\in S_{2}\\ &u_{i}^{*}=-\frac{1}{N}\;\;\forall i:c_{i}>c_{min}+\gamma\\ \end{Bmatrix}.$
Then $\mathrm{conv}\left(\bigcup_{S_{1},S_{2}}U^{*}_{S_{1},S_{2}}\right)$ is equivalently
$\underset{u\in U}{\operatorname*{arg\,min}}\sum_{i=1}^{N}(\frac{1}{N}+u_{i})\cdot c_{i}+\frac{\gamma}{2}\lVert u\rVert_{1}.$ (4)
The theorem explains that the construction of any optimal solution $u^{*}$ essentially reduces to categorizing each of the losses among $\{c_{i}=J(\theta;x_{i},y_{i})\}_{i=1}^{N}$ as “small” or “big”, according to their position in the partitioning of $[c_{min},\infty)=[c_{min},c_{min}+\gamma)\cup[c_{min}+\gamma,\infty)$. For losses that occur at the break points of $c_{min}$ and $c_{min}+\gamma$, this classification can be arbitrary - hence, the use of $S_{1}$ and $S_{2}$ set configurations to capture this degree of freedom.
In particular, those points with losses $c_{i}$ exceeding $c_{min}+\gamma$ are down-weighted to zero and effectively removed from the dataset. And in the event that $c_{max}-c_{min}<\gamma$, no loss reweighting occurs. In this manner, while lasso produces sparse solutions in the model parameter space, RRM produces sparse weight vectors by assigning zero weight to data points with high losses.
It is worth mentioning that upon setting $S_{1}=S_{2}=\emptyset$ in Theorem 3.1, a particular solution for (4) can be derived.
Corollary 3.1.1.
Let $\theta^{*}\in\operatorname*{arg\,min}_{\theta}v(\theta)$ and $\chi=\{i:J(\theta^{*};x_{i},y_{i}))>\min_{j}J(\theta^{*};x_{j},y_{j}))+\gamma\}$. Let $u^{*}$ be given by $\{u^{*}_{i}\}_{i\in{I_{min}}}=\{\frac{|\chi|}{N\cdot|I_{min}|}\}$, $\{u_{i}^{*}\}_{i\in I_{mid}\cup I_{big}}=\{0\}$, and $u_{i}^{*}=\frac{-1}{N}$ for all $i$ such that $c_{i}>c_{min}+\gamma$. Then $(\theta^{*},u^{*})$ solves (3); further, $d_{TV}(p^{N},p^{*})=\frac{|\chi|}{N}.$
In words, transferring $\frac{|\chi|}{N}$ probability mass away from the group of highest-cost examples and distributing it uniformly among the examples comprising $I_{min}$ is optimal.
If $\chi$ converges over the course of any algorithmic scheme, e.g., Algorithm 1, to some set $C$, then we can conclude that these data points are effectively removed from the dataset even if the training of $\theta$ might proceed. This convergence was observed in the experiments of Section 6 (see Table 5). It is hence of possible consideration to tune $\gamma$ for consistency with an estimate $C^{\prime}\in[0,1]$ of labeling error in the dataset $\{(x_{i},y_{i})\}_{i=1}^{N}$. More precisely, we may tune $\gamma$ so that $\frac{|\chi|}{N}\approx C^{\prime}.$